The following set of determination on separate aliquote of a pooled blood serum were reported 103,106,107 and 114meq/L one of the values appears suspect. Should the suspected be retained or discarded at 95% confidence level?






1
Expert's answer
2022-05-23T04:17:03-0400

Let use Q test.


Q=gap/range


gap=114-107=7


range=114-103=11


Q=7/11=0.636


Qcritic=0.829


0.636<0.829


Q<Qcritic

The suspected be retained at 95% confidence level.